# LESS: Large Language Model Enhanced Semi-Supervised Learning for Speech Foundational Models Using in-the-wild Data

### v2

@@ Abstract
Although state-of-the-art Speech Foundation Models can produce high-quality text pseudo-labels, applying Semi-Supervised Learning (SSL) for in-the-wild real-world data remains challenging due to its richer and more complex acoustics compared to curated datasets. To address the challenges, we introduce LESS (Large Language Model Enhanced Semi-supervised Learning), a versatile framework that uses Large Language Models (LLMs) to correct pseudo-labels generated on in-the-wild data. In the LESS framework, pseudo-labeled text from Automatic Speech Recognition (ASR) or Automatic Speech Translation (AST) of the unsupervised data is refined by an LLM, and further improved by a data filtering strategy. Across Mandarin ASR and Spanish-to-English AST evaluations, LESS delivers consistent gains, with an absolute Word Error Rate reduction of 3.8% on WenetSpeech, and BLEU score increase of 0.8 and 0.7, achieving 34.0 on Callhome and 64.7 on Fisher testsets respectively. These results highlight LESS’s effectiveness across diverse languages, tasks, and domains. We have released the recipe as open source to facilitate further research in this area.
@@ 2.1 LESS pipeline
Fig. 1: The proposed LESS pipeline is illustrated taking the ES-to-EN AST task as an example. The process begins by finetuning the SFM at T=0 with supervised Spanish data, as shown by the orange arrow in the diagram. Unsupervised Spanish data collected from real-world sources are processed by a VAD module, which divides the audio into multiple speech segments. These Spanish segments are then transcribed into English using the initial SFM (SFM 0). The resulting English transcriptions are further refined by an LLM. After this, a data filtering step is applied. The filtered, LLM-corrected transcriptions are combined with the original supervised data to further finetune the SFM. This iterative process continues until convergence.
@@ 4.1.2 Discussions
LLM selection. To measure how different LLMs impacts the proposed LESS method, we choose Qwen2.5-coder-32b-instruct 888build.nvidia.com/qwen/qwen2_5-coder-32b-instruct for a comparison with Yi-Large, with the same pseudo-labels from the first iteration of NST in ZH ASR training (Model A in Table 3) and the same prompt. Both LLMs report that they have been trained in Mandarin, making well-suited in our experimental setup. We present results with the filtering threshold 1.0, which means that all the LLM-corrected hypotheses are directly used in NST without data filtering. This allows for a straightforward comparison with the baseline Model A. As shown in Table 3, Yi-Large outperforms Qwen on the wenet_meeting by achieving lower WERs (Model B vs. C). This may be due to the fact that the Qwen model has been post-trained primarily for coding tasks, which could limit its effectiveness in ASR hypothesis correction. In contrast, the Yi-Large model is a more general-purpose LLM, which appears to be better suited for this task. However, a slight performance drop in Model C compared to Model A is observed, possibly because the LLMs do not always refine hypotheses accurately and neither of the two open-source LLMs are post-trained on spoken language data.
@@
Model LLM
Filtering
threshold
WER
Prompting
wenet_meeting

### v3

@@ Abstract
Although state-of-the-art Speech Foundational Models can produce high-quality text pseudo-labels, applying Semi-Supervised Learning (SSL) for in-the-wild real-world data remains challenging due to its richer and more complex acoustics compared to curated datasets. To address the challenges, we introduce LESS (Large Language Model Enhanced Semi-supervised Learning), a versatile framework that uses Large Language Models (LLMs) to correct pseudo-labels generated on in-the-wild data. In the LESS framework, pseudo-labeled text from Automatic Speech Recognition (ASR) or Automatic Speech Translation (AST) of the unsupervised data is refined by an LLM, and further improved by a data filtering strategy. Across Mandarin ASR and Spanish-to-English AST evaluations, LESS delivers consistent gains, with an absolute Word Error Rate reduction of 3.8% on WenetSpeech, and BLEU score increase of 0.8 and 0.7, achieving 34.0 on Callhome and 64.7 on Fisher testsets respectively. These results highlight LESS’s effectiveness across diverse languages, tasks, and domains. We have released the recipe as open source to facilitate further research in this area.
@@ 2.1 LESS pipeline
Fig. 1: The proposed LESS pipeline is illustrated taking the ES-to-EN AST task as an example. The process begins by finetuning the SFM at T=0 with supervised Spanish data (orange arrow). Unsupervised Spanish data collected from real-world sources are segmented by a VAD module. These Spanish segments are then transcribed into English using the initial SFM (SFM 0). The resulting English transcriptions are further refined by an LLM. After this, a data filtering step is applied. The filtered, LLM-corrected transcriptions are combined with the original supervised data to further finetune the SFM. This iterative process continues until convergence.
@@ 4.1.2 Discussions
LLM selection. To measure how different LLMs impacts the proposed LESS method, we choose Qwen2.5-coder-32b-instruct 888build.nvidia.com/qwen/qwen2_5-coder-32b-instruct for a comparison with Yi-Large, with the same pseudo-labels from the first iteration of NST in ZH ASR training (Model A in Table 3) and the same prompt. The Yi-Large model is trained with 3.1 trillion tokens of English and Chinese corpora and the training data of Qwen2.5-coder-32b-instruct includes 5.5 trillion tokens from source code, text-code grounding and synthetic data. Both LLMs report that they have been trained in Mandarin, making them well-suited in our experimental setup. We present results with the filtering threshold 1.0, which means that all the LLM-corrected hypotheses are directly used in NST without data filtering. This allows for a straightforward comparison with the baseline Model A. As shown in Table 3, Yi-Large outperforms Qwen on the wenet_meeting by achieving lower WERs (Model B vs. C). This may be due to the fact that the Qwen model has been post-trained primarily for coding tasks, which could limit its effectiveness in ASR hypothesis correction. In contrast, the Yi-Large model is a more general-purpose LLM, which appears to be better suited for this task. However, a slight performance drop in Model C compared to Model A is observed, possibly because the LLMs do not always refine hypotheses accurately and neither of the two open-source LLMs are post-trained on spoken language data.
@@
Model LLM Filtering threshold WER Prompting wenet_meeting